a small plane cruises at 180 mph. it takes the same amount of time to travel 7 miles with the wind as it takes to travel 5 miles against the wind. what is the wind speed?
Answer by dabanfield(803)

a small plane cruises at 180 mph. it takes the same amount of time to travel 7 miles with the wind as it takes to travel 5 miles against the wind. what is the wind speed?
Let w be the wind speed and t the time traveled. Then using distance = rate*time we have:
1.) 7 = (180+w)*t (with the wind)
2.) 5 = (180-w)*t (against the wind)
Expanding we have:
From 1.) we have t = 7/(180+w)
Substituting 7/(180+w) for t in 2.) we have:
5 = (180-w)*(7/(180+w))
Multiply both sides above by (180+w):
5*(180+w) = (180-w)*7
900+5w = 1260-7w
Solve the above for w.